OpenStudy (anonymous):

Simplify sqrt(2)(x - sqrt(5))?

OpenStudy (anonymous):

\[\sqrt{2}(x - \sqrt{5})\]

OpenStudy (anonymous):

And then the possible answers are.. \[\sqrt{2x} - \sqrt{10}\] \[10\sqrt{2x}\] \[x \sqrt{2} - \sqrt{10}\] \[2x \sqrt{2}\]

OpenStudy (anonymous):

Help would be muchly appreciated. :o

OpenStudy (anonymous):

but I think it might be \[\sqrt{2x} - \sqrt{10}\] Is that right? D: